# VFScale: Intrinsic Reasoning through Verifier-Free Test-time Scalable Diffusion Model

### v3

@@ 4.3 Inference of VFScale
To fully harness the potential of the diffusion model with the higher-efficiency search algorithm, we propose a novel hybrid Monte Carlo Tree Search denoising (hMCTS denoising) method, which progressively applies best-of-$N$ (BoN) and Monte Carlo Tree Search (MCTS) denoising in sequence, as detailed below, throughout the denoising process. As demonstrated in Algorithm  1, we employ BoN for the diffusion process in the early diffusion stages, which introduces $L$ initial noises to maintain a consistent number of function evaluations (NFE) per example during the diffusion process. [^7] In this paper, we primarily report $K=N_{r}$ MCTS denoising, where ensuring $K=N_{r},L=N_{r}$ guarantees that the NFE (Number of Function Evaluations) for each case of BoN, MCTS denoising, and hMCTS denoising remains the same. Subsequently, hMCTS denoising utilizes the MCTS denoising [^8] For a detailed explanation of the distinctions between hMCTS denoising, MCTS denoising, and BoN, as well as the impact of MCTS denoising start step $t_{s}$, please refer to Appendix D. to iteratively perform the denoising process until the termination state ${\bm{x}}_{0}$ is reached. This approach enables MCTS to more accurately estimate node value when the noise is relatively small, thereby preventing the premature exclusion of potentially promising nodes. Next, we will detail the MCTS denoise process.
We treat the current diffusion state ${\bm{x}}_{t}$ as the state, the noise to remove as the action, and model the denoising process of the diffusion model as a Markov Decision Process (MDP). Therefore, a node in MCTS represents the state ${\bm{x}}_{t}$, along with its current visit count $N({\bm{x}}_{t})$ and state value $Q({\bm{x}}_{t})$. A terminal node is defined as a node whose denoising step is 0. In this context, we use $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ and $E_{\theta}(\bm{x}_{t},t)$ of the energy-based diffusion as the policy network and value network, respectively. Each deepening of the search tree corresponds to a single denoising step. Similar to MCTS in AlphaGo silver2016mastering and AlphaZero  silver2017mastering , each rollout in MCTS consists of four core steps: selection, expansion, simulation, and backpropagation, as illustrated in Fig. 2 and Algorithm  1 in Appendix A:
(1) Selection: Based on the Upper Confidence Bound (UCB) from AlphaGo silver2016mastering , starting from the current root node state $\bm{x}_{t}$, we select a child node based on the following adjusted UCB of MCTS-enhanced diffusion formula until a leaf node $\{{\bm{x}}_{t^{\prime}},Q({\bm{x}}_{t^{\prime}}),N({\bm{x}}_{t^{\prime}})\}$ is reached:
@@
where $Q(\bm{x}_{t},\bm{a}_{t})$ represent the value of children node, action $\bm{a}_{t}$ includes predicted noise $\bm{\epsilon}_{\theta}$ and random Gaussian noise $\bm{\epsilon}$, $n_{i}$ represents the number of visits to node $i$, $N_{i}$ represents the number of visits to the parent node of $i$, and $c$ is the exploration hyperparameter.
(2) Expansion: Unless the state of the node reached is a terminal state $\bm{x}_{0}$, we expand the children of the selected node by choosing an action and creating new nodes based on the action. For the expansion step of MCTS denoising, we perform a denoising step and add different but limited Gaussian noise. This results in $K$ distinct branches $\left\{\bm{x}_{t^{\prime}-1}^{(k)}\mid k=0,1,\cdots,K-1\right\}$, where each child node $\bm{x}_{t^{\prime}-1}^{(k)}$ is derived as following equation adjusted from Eq. 3 :
@@
(3) Simulation: We randomly select a child node and perform a random simulation of the MDP until reaching a terminal state. For the simulation of MCTS denoising, we use DDIMs song2020denoising for fast sampling to obtain $\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})})$ from the randomly chosen child node state ${\bm{x}}_{t^{\prime}-1}^{(k^{*})}$, and then use $E_{\theta}(\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})}))$ as the reward for backpropagation instead of reward from external verifier. [^9] Although our method is based on an energy-based diffusion model, the proposed hMCTS is also applicable to conventional noise-predicting diffusion models, albeit requiring an external verifier. Additional results are provided in the Appendix E.
@@
After $N_{r}$ rollouts, we select $\bm{x}_{t-1}^{(k)}$ with the largest value $\frac{Q({\bm{x}}_{t-1}^{(k)})}{N({\bm{x}}_{t-1}^{(k)})}$ as the state $\bm{x}_{t-1}$ for the next denoising step. The next MCTS denoising process starts from this state and proceeds until the termination state $\bm{x}_{0}$. Here, the depth of the search tree is decided by three elements: the number of rollout steps $N_{r}$, the maximum number of branches $K$ for each node, and the search policy.
To summarize, our proposed paradigm, VFScale, enhances performance-energy consistency and ease of sampling through LRNCL loss and KL regularization, respectively (Sec. 4.2), while fully unlocking the test-time scalability of diffusion models via hMCTS denoising (Sec. 4.3).
@@ 5 Experiments
To answer these questions, we conduct extensive experiments on two challenging reasoning tasks: Maze and Sudoku. In Section  5.1, we find that the naïvely trained energy-based diffusion model du2024learning falls short of scaling up during test time and identify the key reason as the poor quality of the energy landscape, particularly its lack of performance-energy consistency. Addressing this In Section 5.2, the training methods of VFScale demonstrate a significant impact on unbinding the scalability of diffusion models. As shown in Fig. 1, with the model trained with VFScale training methods, hMCTS denoising can better release the scalability, enabling generalization to much more challenging tasks during inference (Sec. 5.3).
In Maze experiments, the datasets are generated by ivanitskiy2023configurable , and the diffusion model is mainly trained with Maze sizes of up to $6\times 6$ while tested on harder datasets with sizes significantly larger than $6\times 6$ (Fig. 1). For Sudoku experiments, the basic setting is adopted from du2024learning where the diffusion model is trained on SAT-Net dataset with 31 to 42 given entries wang2019satnet and tested on the harder RRN dataset with 17 to 34 given entries palm2018recurrent with fewer given entries. All models and inference methods are evaluated with solving success rate. Here successful solving means the predicted solution exactly matches the ground-truth solution on all entries, a very stringent metric [^10] For example, a successful solving of a $15\times 15$ Maze needs to predict all $31\times 31=961$ entries correctly (for each grid cell, predict path/not path/wall), where the path length from start to target is on the order of $10^{2}$.. More experimental details can be found in Appendix B, Appendix C, and Appendix H, respectively.
All the training methods are compared with the original training pipeline in du2024learning . LRNCL and KL represent the loss terms ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, respectively in Eq. 6. VFScale tr. w/o LRNCL, VFScale tr. w/o KL, and VFScale tr. (ours) represent the three training methods of VFScale, the last one being the full version. The MCTS denoising and hMCTS denoising are compared with BoN with the same computational budget. The experiment code can be found at the repo.
@@ 5.1 Test-Time Scaling bottleneck of energy-based diffusion models
Test-time scaling up the diffusion model to solve harder reasoning problems is challenging. When tested on more challenging tasks, diffusion models trained with the original method fail to achieve substantial gains, even under larger inference budgets. Specifically, as shown in Table 1, models trained on simpler tasks with original training methods exhibit significantly poorer performance on more complex tasks. Success rate plunges from 100% on Maze size $6\times 6$ to 6% on $15\times 15$ , and from 32% on Sudoku with 33 given entries to below 5% at 25 and 0% at 21 givens. Building upon this, we conduct best-of-N to examine whether test-time scaling can enhance the performance of the originally trained model. Table 3’s “Original, BoN” row reveals that enlarging the inference budget $N$ offers marginal gains—Maze success climbs by at most 5%, while Sudoku remains below 30% even at $N=320$. These indicate that naïve test-time scaling of diffusion models fails to deliver meaningful improvements. Next, we analyze the underlying causes from the perspectives of training and inference, respectively.
Table 1: Success rate across different grid sizes of Maze and various number of given entries for naïve inference ($N=1$) for comparison of the generalization ability of models obtained by different training methods. Let $M$ denote the grid size of Maze and $D$ denote the number of given digits in Sudoku. Original denotes the original training method in du2024learning . VFScale tr. (ours) denotes the training method of VFScale, and all subsequent descriptions will follow this definition. Bold font denotes the best model, and an underline denotes the second-best model. The same markings are used in the tables below.
@@
VFScale tr. w/o LRNCL, BoN (Ground Truth) 0.2500 0.5078 0.5938 0.6562 0.7109 0.7422 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219
First, there is still considerable room for improvement on the training side to fully unlock the test-time scalability of diffusion models. As shown in Table 2, even when ground truth is used to guide test-time scaling, the success rate on the Maze task increases by only 7% over the energy-guided scaling method at an inference budget of $N=161$, remaining below 20%. On the Sudoku task, raising $N$ to 321 yields under a 2% improvement compared with energy-guided scaling, still below 30%. These findings indicate that employing the learned energy function as the verifier can lead to misestimation; the performance–energy consistency must be further improved. Moreover, Table 4 reports that the originally trained model achieves only about 70% performance–energy consistency in quantitative evaluation, underscoring a significant mismatch between the energy-based verifier and actual performance. In short, these results highlight the urgent need to refine training methods to improve the quality of energy landscape.
Second, substantial advancements in test-time scaling methodologies are still required to fully harness the inherent scalability of diffusion models. As shown in Table 2’s first and second rows, regardless of whether ground truth is as a verifier, the performance of BoN quickly plateaus. On the Maze task, increasing the inference budget $N$ from 1 to 161 raises the success rate from 6% to 17%—a gain of only 10%—and the rate of improvement slows markedly; a similar trend is observed on Sudoku, where success rate growth decelerates significantly and the maximum success rate remains below 30%. These findings suggest that there remains substantial room to enhance the efficiency of test-time scaling methods.
@@ 5.2 The effect of VFScale training methods
Table 4: Performance-energy consistency of BoN on Maze with grid size 15 to test the effect of LRNCL loss. Here, $L=N$. Details of consistency calculation can be found in Appendix A.
@@
In this subsection, we evaluate the effectiveness of the LRNCL and KL losses in VFScale compared with the baseline model trained without these losses. First, as shown in Table 3’s “N=1” columns, even without test-time scaling, integrating either the LRNCL loss or the KL loss produces significant improvements: on the Maze task, the success rate rises from 6% to 28%, and on Sudoku it likewise increases from 9% to 16%. Second, Table 4 also shows that adding the LRNCL loss alone produces a steady improvement in performance–energy consistency of over 10%. Finally, when we apply BoN for test-time scaling of the diffusion model, Table 3 demonstrates that using only the LRNCL loss, only the KL loss, or both, boosts the Maze success rate by more than 60% and the Sudoku success rate by over 10%. Meanwhile, as illustrated in the Fig. 4, after incorporating the LRNCL loss and the KL loss, the solutions produced by the model at different denoising steps are all noticeably closer to the ground truth solutions. Together, these findings provide strong evidence that the LRNCL and KL losses significantly enhance the test-time scalability of diffusion models.
@@ 5.3 Test-time Scalability of VFScale
Using the diffusion model trained with VFScale training methods, we evaluate various inference approaches to validate the efficacy of our inference methods described in Section 4.3. As shown in Table 5, in the Maze experiment, the MCTS denoising method slightly outperforms BoN, while our hMCTS denoising yields a significantly higher success rate, with a maximum improvement of approximately 18% than BoN. Moreover, as the budget $N$ increases, the performance gap between hMCTS denoising and BoN widens. In the Sudoku experiment, hMCTS denoising also consistently outperforms BoN, with a maximum improvement of 5%. As illustrated in the scaling curve in Fig. 3, hMCTS denoising with the model trained with additional LRNCL loss and KL loss shows a marked improvement compared to both other inference methods. At the same time, the rate of improvement for other inference methods clearly slows down compared to hMCTS denoising. These results provide strong evidence that our inference method can effectively scale up during test time, offering a clear advantage over BoN.
@@ 6 Conclusion
In this work, we have introduced the Verifier-free Test-time Scalable Diffusion Model (VFScale), a novel framework to achieve scalable intrinsic reasoning to tackle more complex reasoning tasks than in training. VFScale explores applying the varying number-of-sample scaling method with learned energy as a verifier. Concretely, faced with a low-quality energy landscape and lack of efficient search algorithms, VFScale is composed of two innovations on training and inference to address these challenges. On the training side, VFScale introduces two auxiliary training losses, ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, to improve the energy landscape. This, in turn, ensures that using the learned energy function as a verifier is effective for test-time scaling. On the inference side, VFScale integrates hybrid Monte Carlo Tree Search (hMCTS) denoising to better leverage the model’s test-time scalability. We have conducted extensive experiments on Maze and Sudoku to validate the efficacy of VFScale and believe that VFScale offers a robust solution for test-time scaling, unlocking the notable potential of diffusion models for complex reasoning. Limitations and future directions are discussed in Appendix G.
Appendix A Related algorithms and metric caculation

### v4

@@ 4.3 Inference of VFScale
To fully exploit the potential of the diffusion model with a more efficient search algorithm, we propose a novel hybrid Monte Carlo Tree Search denoising (hMCTS denoising) method, which progressively applies best-of-$N$ (BoN) and Monte Carlo Tree Search (MCTS) denoising in sequence, as detailed below, throughout the denoising process. As demonstrated in Algorithm  1, we employ BoN for the diffusion process in the early diffusion stages, which introduces $L$ initial noises to maintain a consistent number of function evaluations (NFE) per example during the diffusion process. [^6] In this paper, we primarily report $K=N_{r}$ MCTS denoising, where ensuring $K=N_{r},L=N_{r}$ guarantees that the NFE (Number of Function Evaluations) for each case of BoN, MCTS denoising, and hMCTS denoising remains the same. Allocating an identical NFE is the basis for a fair comparison of the search strategies’ effectiveness, as detailed in Appendix F. Subsequently, hMCTS denoising utilizes the MCTS denoising [^7] For a detailed explanation of the distinctions between hMCTS denoising, MCTS denoising, and BoN, as well as the impact of MCTS denoising start step $t_{s}$, please refer to Appendix D. to iteratively perform the denoising process until the termination state ${\bm{x}}_{0}$ is reached. This approach enables MCTS to more accurately estimate node values when the noise is relatively small, thereby preventing the premature exclusion of potentially promising nodes. Next, we will detail the MCTS denoise process.
We treat the current diffusion state ${\bm{x}}_{t}$ as the state, the noise to remove as the action, and model the denoising process of the diffusion model as a Markov Decision Process (MDP). Therefore, a node in MCTS represents the state ${\bm{x}}_{t}$, along with its current visit count $N({\bm{x}}_{t})$ and the state value $Q({\bm{x}}_{t})$. A terminal node is defined as a node whose denoising step is 0. In this context, we use $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ and $E_{\theta}(\bm{x}_{t},t)$ of the energy-based diffusion as the policy network and value network, respectively. Each deepening of the search tree corresponds to a single denoising step. Similar to MCTS in AlphaGo (Silver et al., 2016) and AlphaZero  (Silver et al., 2017), each MCTS rollout consists of four core steps: selection, expansion, simulation, and backpropagation, as illustrated in Fig. 2 and Algorithm  1 in Appendix A:
(1) Selection: Based on the Upper Confidence Bound (UCB) from AlphaGo (Silver et al., 2016), starting from the current root node state $\bm{x}_{t}$, we select a child node based on the following adjusted UCB of MCTS-enhanced diffusion formula until a leaf node $\{{\bm{x}}_{t^{\prime}},Q({\bm{x}}_{t^{\prime}}),N({\bm{x}}_{t^{\prime}})\}$ is reached:
@@
where $Q(\bm{x}_{t},\bm{a}_{t})$ represents the value of children node, action $\bm{a}_{t}$ includes predicted noise $\bm{\epsilon}_{\theta}$ and random Gaussian noise $\bm{\epsilon}$, $n_{i}$ represents the number of visits to node $i$, $N_{i}$ represents the number of visits to the parent node of $i$, and $c$ is the exploration hyperparameter.
(2) Expansion: Unless the state of the node reached is a terminal state $\bm{x}_{0}$, we expand the children of the selected node by choosing an action and creating new nodes based on the action. For the expansion step of MCTS denoising, we perform a denoising step and add different but limited Gaussian noise. This results in $K$ distinct branches $\left\{\bm{x}_{t^{\prime}-1}^{(k)}\mid k=0,1,\cdots,K-1\right\}$, where each child node $\bm{x}_{t^{\prime}-1}^{(k)}$ is derived as the following equation adjusted from Eq. 3 :
@@
(3) Simulation: We randomly select a child node and perform a random simulation of the MDP until a terminal state is reached. For the simulation of MCTS denoising, we use DDIM (Song et al., 2021a) for fast sampling to obtain $\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})})$ from the randomly chosen child node state ${\bm{x}}_{t^{\prime}-1}^{(k^{*})}$, and then use $E_{\theta}(\hat{\bm{x}}_{0}(\bm{x}_{t^{\prime}-1}^{(k^{*})}))$ as the reward for backpropagation instead of a reward from external verifier. [^8] Although our method is based on an energy-based diffusion model, the proposed hMCTS is also applicable to conventional noise-predicting diffusion models, albeit requiring an external verifier. Additional results are provided in Appendix E.
@@
After $N_{r}$ rollouts, we select $\bm{x}_{t-1}^{(k)}$ with the highest value $\frac{Q({\bm{x}}_{t-1}^{(k)})}{N({\bm{x}}_{t-1}^{(k)})}$ as the state $\bm{x}_{t-1}$ for the next denoising step. The next MCTS denoising process starts from this state and proceeds to the termination state $\bm{x}_{0}$. Here, the depth of the search tree is decided by three elements: the number of rollout steps $N_{r}$, the maximum number of branches $K$ for each node, and the search policy.
In summary, our proposed paradigm, VFScale, enhances performance-energy consistency and ease of sampling through MRNCL loss and KL regularization, respectively (Sec. 4.2), while fully unlocking the test-time scalability of diffusion models via hMCTS denoising (Sec. 4.3).
@@ 5 Experiments
To answer these questions, we conduct extensive experiments on two challenging reasoning tasks: Maze and Sudoku, evaluating them in out-of-distribution (OOD) settings where test instances are substantially more complex than the training data and cause naive methods to fail. [^9] More experimental details can be found in Appendix B, Appendix C, and Appendix J, respectively.. In Section  5.1, we find that the naïvely trained energy-based diffusion model (Du et al., 2024) falls short of scaling up during test time and identify the key reason as the poor quality of the energy landscape, particularly its lack of performance-energy consistency. Addressing this In Section 5.2, the training methods of VFScale demonstrate a significant impact on unbinding the scalability of diffusion models. As shown in Fig. 1, with the model trained with VFScale training methods, hMCTS denoising can better release the scalability, enabling generalization to much more challenging tasks during inference (Sec. 5.3).
@@ 5.1 Test-Time Scaling bottleneck of energy-based diffusion models
Test-time scaling up the diffusion model to solve harder reasoning problems is challenging. When tested on more challenging tasks, diffusion models trained with the original method fail to achieve substantial gains, even under larger inference budgets. Specifically, as shown in Table 1, models trained on simpler tasks with original training methods exhibit significantly poorer performance on more complex tasks. Success rate plunges from 100% on Maze size $6\times 6$ to 6% on $15\times 15$ , and from 32% on Sudoku with 33 given entries to below 5% at 25 and 0% at 21 givens. Building upon this, we conduct best-of-N to examine whether test-time scaling can enhance the performance of the originally trained model. Table 3’s “Original, BoN” row reveals that enlarging the inference budget $N$ offers marginal gains—Maze success rises by at most 5%, while Sudoku remains below 30% even at $N=320$. These indicate that naïve test‑time scaling of diffusion models fails to deliver meaningful improvements. Next, we analyze the underlying causes from the perspectives of training and inference, respectively.
Table 1: Success rate across different grid sizes of Maze and various number of given entries for naïve inference ($N=1$) for comparison of the generalization ability of models obtained by different training methods. Let $M$ denote the grid size of Maze and $D$ denote the number of given digits in Sudoku. Original denotes the original training method in Du et al. (2024). VFScale tr. (ours) denotes our full training method, and this naming convention is used for following figures and tables. Bold font denotes the best model, and an underline denotes the second-best model. The same markings are used in the tables below.
@@
VFScale tr. w/o MRNCL, BoN (Ground Truth) 0.2500 0.5078 0.5938 0.6562 0.7109 0.7422 0.1094 0.2578 0.2969 0.3438 0.3750 0.3828 0.4219
First, there is still considerable room for improvement on the training side to fully unlock the test-time scalability of diffusion models [^10] As shown in the Appendix D.1, training an effective external verifier is non-trivial and a dense reward signal is crucial, which underscores the value of using the learned energy as an intrinsic verifier.. As shown in Table 2, even when ground truth is used to guide test-time scaling, the success rate on the Maze task increases by only 7% over the energy-guided scaling method at an inference budget of $N=161$, remaining below 20%. In the Sudoku task, raising $N$ to 321 yields only a 2% improvement compared with energy-guided scaling, still below 30%. These findings indicate that employing the learned energy function as the verifier can lead to misestimation; the performance–energy consistency must be further improved. Moreover, Table 4 reports that the originally trained model achieves only about 70% performance–energy consistency in quantitative evaluation, underscoring a significant mismatch between the energy-based verifier and actual performance. In short, these results highlight the urgent need to refine training methods to improve the quality of the energy landscape.
Second, substantial advancements in test-time scaling methodologies are still required to fully harness the inherent scalability of diffusion models. As shown in Table 2’s first and second rows regardless of whether ground truth is as a verifier, the performance of BoN quickly plateaus. In the Maze task, increasing the inference budget $N$ from 1 to 161 raises the success rate from 6% to 17%—a gain of only 10% — and the rate of improvement slows markedly; a similar trend is observed in Sudoku, where success rate growth decelerates significantly and the maximum success rate remains below 30%. These findings suggest that there remains substantial room to improve the efficiency of test-time scaling methods.
@@ 5.2 The effect of VFScale training methods
Table 4: Performance-energy consistency of BoN on Maze with grid size 15 to test the effect of MRNCL loss. Here, $L=N$. Details of consistency calculation can be found in Appendix A.
@@
In this subsection, we evaluate the effectiveness of the MRNCL and KL losses in VFScale compared with the baseline model trained without these losses. First, as shown in Table 3’s “N=1” columns, even without test-time scaling, integrating either the MRNCL loss or the KL loss produces significant improvements: on the Maze task, the success rate rises from 6% to 28%, and on Sudoku it likewise increases from 9% to 16%. Second, Table 4 also shows that adding the MRNCL loss alone produces a steady improvement in performance–energy consistency of over 10%. Finally, when we apply BoN for test-time scaling of the diffusion model, Table 3 demonstrates that using only the MRNCL loss, only the KL loss, or both, boosts the Maze success rate by more than 60% and the Sudoku success rate by over 10%. Meanwhile, as illustrated in the Fig. 4, after incorporating the MRNCL loss and the KL loss, the solutions produced by the model at different denoising steps are all noticeably closer to the ground truth solutions. Together, these findings provide strong evidence that the MRNCL and KL losses significantly enhance the test-time scalability of diffusion models.
@@
Trade-off between Naïve Inference and Scalability. A critical observation from Table 1 is that the full VFScale model occasionally underperforms ablated variants (e.g., "w/o KL") in the naïve $N=1$ setting. This is a deliberate trade-off central to our contribution. The ablated models produce “sharper” energy landscapes that may yield a confident local minimum in a single step (hence higher $N=1$ scores) but are often riddled with poor local optima that trap search algorithms. In contrast, our joint objectives ($\mathcal{L}_{MRNCL}+\mathcal{L}_{KL}$) are explicitly designed to smooth the landscape, intentionally sacrificing $N=1$ sharpness to ensure global navigability. The payoff for this sacrifice is superior scalability. For instance, in the Maze task (Table 3), the "w/o KL" model starts strong at 0.2812 but plateaus at 0.6250 ($N=161$). Conversely, our full model starts lower (0.2656) but scales dramatically to 0.7031 (BoN) and further to 0.8828 with hMCTS (Table 5). Thus, the lower $N=1$ performance is a symptom of a successful optimization for test-time search potential.
@@ 5.3 Test-time Scalability of VFScale
Using the diffusion model trained with VFScale training methods, we evaluate various inference approaches to validate the efficacy of our inference methods described in Section 4.3. As shown in Table 5, in the Maze experiment, the MCTS denoising method slightly outperforms BoN, while our hMCTS denoising yields a significantly higher success rate, with a maximum improvement of approximately 18% than BoN. Moreover, as budget $N$ increases, the performance gap between hMCTS denoising and BoN widens. In the Sudoku experiment, hMCTS denoising also consistently outperforms BoN, with a maximum improvement of 5%. As illustrated in the scaling curve in Fig. 3, hMCTS denoising with the model trained with additional MRNCL loss and KL loss shows a marked improvement compared to both other inference methods. At the same time, the rate of improvement for other inference methods clearly slows down compared to hMCTS denoising. These results provide strong evidence that our inference method can effectively scale up during test time, offering a clear advantage over BoN.
@@ 6 Conclusion
In this work, we have introduced the Verifier-free Test-time Scalable Diffusion Model (VFScale), a novel framework to achieve scalable intrinsic reasoning to tackle more complex reasoning tasks than in training. VFScale explores applying the varying number-of-sample scaling method with learned energy as a verifier. Concretely, faced with a low-quality energy landscape and the lack of efficient search algorithms, VFScale is composed of two innovations in training and inference to address these challenges. On the training side, VFScale introduces two auxiliary training losses, ${\mathcal{L}}_{\text{MRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, to improve the energy landscape. This, in turn, ensures that the learned energy function is an effective verifier for test-time scaling. On the inference side, VFScale integrates hybrid Monte Carlo Tree Search (hMCTS) denoising to better leverage the model’s test-time scalability. We have conducted extensive experiments on Maze and Sudoku to validate the efficacy of VFScale and believe that VFScale offers a robust solution for test-time scaling, unlocking the notable potential of diffusion models for complex reasoning. Limitations and future directions are discussed in Appendix I.
Acknowledgments
This work was supported by the Westlake University Center for High-performance Computing. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies or endorsements of the funding entities.
Appendix A Related algorithms and metric calculation